Problem 9: A ball is thrown straight up at time t = 0 with an initial speed of 15 m/s. Take the point of release to be y0 = 0 and upwards to be the positive direction. No Attempt No Attempt 13% Part

(a) Calculate the displacement at the time of 0.50 s. No Attempt No Attempt 13% Part

(b) Calculate the velocity at the time of 0.50 s. No Attempt No Attempt 13% Part

(c) Calculate the displacement at the time of 1.0 s. No Attempt No Attempt 13% Part

(d) Calculate the velocity at the time of 1.0 s. No Attempt No Attempt 13% Part

(e) Calculate the displacement at the time of 1.5 s. No Attempt No Attempt 13% Part

(f) Calculate the velocity at the time of 1.5 s. No Attempt No Attempt 13% Part

(g) Calculate the displacement at the time of 2.0 s. show answer No Attempt 13% Part

(h) Calculate the velocity at the time of 2.0 s.
Answer #1

Date: 11 Page No. Ans. given u=15ms throum upwards, so g=-9.8 m/s2 Can t = 0. 505 S = 1 g ² + ut S = (15)(0:59x1 698960592 SDate: 1 Page No. distance travelled at t=1.08 = displacement displacement = 11.1 m D) Velocity at t- tiosec V = ut gt V = 15+Date: / Page No. displacement at t=2 see on o. We know that v=0 m/s, u=15 m/s, g = 9.8m. its time taken by ball to reach at mDate: 11 Page No. Velocity at t=asec. U=0,9 = 98 mis? v=utgt v=o+(9.8 (0.47) 1 = 4.606 m/s So, at t= 28ee relouty will be 4.